OpenStudy (nuccioreggie):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $ Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in by

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (anonymous):

64

OpenStudy (nuccioreggie):

Whats the answer for part A

OpenStudy (michele_laino):

if I call with p the price in September, then we have this equation: \[\Large p + \frac{{20}}{{100}}p = 1.50\] solve for p, please

OpenStudy (nuccioreggie):

420

OpenStudy (nuccioreggie):

Whats the answer for part A my friend

OpenStudy (michele_laino):

hint: next steps are: \[\begin{gathered} p\left( {1 + 0.2} \right) = 1.5 \hfill \\ p = \frac{{1.5}}{{1.2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (nuccioreggie):

7.5?

OpenStudy (nuccioreggie):

Whats the CORECT ANSWER!

OpenStudy (michele_laino):

what is: 1.5:1.2=...?

OpenStudy (nuccioreggie):

1.8

OpenStudy (michele_laino):

no it is 1.25